# NuwaTS: a Foundation Model Mending Every Incomplete Time Series

### v2

@@ 3.3 Domain-Specific Fine-Tuning
To achieve a domain-specific model, we borrow the idea of P-tuningv2 strategy [17]. Unlike pre-training phase, where the domain-specific embedding $\mathbf{k}$ is only added to the input embedding. We incorporate this domain-specific information into every layer of the PLM. Initially, we employ a domain-transfer layer—a two-layer MLP network—to map $\mathbf{k}\in\mathbb{R}^{D}$ to $\boldsymbol{\mathcal{\hat{K}}}\in\mathbb{R}^{2\times\text{Layer}\times D}$. Following the P-tuningv2 approach, we also randomly initialize a continuous prompt $\boldsymbol{\mathcal{P}}\in\mathbb{R}^{2\times\text{Layer}\times D}$ for each layer. We then combine the randomly initialized $\boldsymbol{\mathcal{P}}$ with the domain-specific information $\boldsymbol{\mathcal{\hat{K}}}$ to serve as the domain-specific prefix Key and Value in the PLM’s hidden layers. Thus, we have $[\mathbf{Key}_{p},\mathbf{Value}_{p}]=\boldsymbol{\mathcal{P}}+\beta\boldsymbol{\mathcal{\hat{K}}}$, where $\beta=0.01$.
@@
Here, $\mathbf{Key}_{p}$ and $\mathbf{Value}_{p}\in\mathbb{R}^{\text{Layer}\times D}$ contain the domain-specific prefix key and value for each layer and participate in the subsequent attention computation. During fine-tuning on time series data from the target domain, we freeze all parameters except for the randomly initialized $\boldsymbol{\mathcal{P}}$ and the domain-transfer layer. The fine-tuning process is illustrated in Figure 4. If these two trainable modules are removed, our domain-specific model reverts to the original one-for-all foundation model. Therefore, the fine-tuning strategy we propose is essentially “plug-and-play”.
@@ 4 Experiment
We conducted comprehensive experiments on NuwaTS across ten commonly used datasets from various domains. To facilitate comparison, we provided three versions of the model based on the GPT2 architecture: a specific model trained on a single domain, an one-for-all model trained on a general fused dataset (17.6M collected samples from various domains), and a fine-tuned model for a specific domain (fine-tuned based on the one-for-all model). Moreover, we evaluated BERT [7] and LLaMA2 [35] as backbones for NuwaTS (details in Appendix D.3). To assess its zero-shot capability, we trained another GPT2-based model on a transportation domain dataset, LargeST [19] (comprising 100.1 million collected samples) and evaluate it on time series from other domains.
@@
Baselines. Following the configuration outlined in [8], we evaluated two naive imputation methods: Median, where missing values are filled with the median value, and Last, where missing values are filled with the last previous observations. To ensure a fair comparison, all methods followed a unified pipeline [^2] https://github.com/thuml/Time-Series-Library. We assessed two classic deep learning-based imputation models, BRITS[6] and SAITS[8]. We also compared our model with other foundation models for time series such as DLinear[41], PatchTST[26], iTransformer[20], TimesNet[38], Autoformer[39], Fedformer[43], and GPT4TS[44]. We trained GPT4TS separately on single datasets with the number of layers set to 6 to align with NuwaTS. We trained all these models using a domain-specific approach. Since PatchTST is a transformer-based model with bi-directional attention, we also trained and tested it on the general fused dataset.
@@
Setups. We partitioned the dataset along the sensor (variable) dimension into training, validation, and test sets in a 1:1:1 ratio. This division simulates the process of collecting relatively complete time series data from a few sensors or variables with lower missing rates, training a model with this data, and then using it to impute data from other sensors or variables that have higher missing rates. The input time series length $L$ is 96, and we trained all baselines under random missing rates (sampled from 0.1 to 0.9) and tested them separately under 9 missing rates: 0.1, 0.2,…, 0.9. We used a total of ten datasets for domain-specific training and mixed them into a general fused dataset for one-for-all training. Appendix A.1 shows the details of datasets. For zero-shot training, we chose LargeST [^3] https://github.com/liuxu77/LargeST, a large-scale traffic dataset including a total of 8,600 time series and over 100 million samples, to train NuwaTS and conduct zero-shot experiments on other domains such as ETTs, weather and electricity.
@@
Table 1: Imputation results are presented for ETTs, ECL, and weather datasets with nine test missing rates: 0.1, 0.2, …, 0.9. All results are averaged across the nine different missing rates. A lower MSE or MAE signifies better imputation performance. Red highlights the best results, while Blue indicates the second-best.
[b] Model ETTh1 ETTh2 ETTm1 ETTm2 ECL Weather MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Median 0.723 0.611 0.728 0.472 0.699 0.584 0.744 0.464 1.010 0.834 0.998 0.497 Last 0.432 0.476 0.089 0.149 0.336 0.413 0.059 0.122 0.965 0.826 0.731 0.349 Autoformer(2021) 0.552 0.547 0.472 0.420 0.346 0.409 0.211 0.312 0.137 0.262 0.610 0.450 Fedformer(2022) 0.354 0.436 0.538 0.429 0.076 0.185 0.055 0.156 0.129 0.258 0.237 0.211 Dlinear(2023) 0.356 0.414 0.245 0.301 0.274 0.357 0.266 0.351 0.317 0.419 0.355 0.309 iTransformer(2024) 0.639 0.572 0.369 0.376 0.352 0.413 0.356 0.416 0.092 0.200 0.515 0.412 BRITS(2018)1 0.213 0.313 0.117 0.172 0.096 0.183 0.071 0.123 0.317 0.427 0.793 0.474 TimesNet(2022) 0.166 0.280 0.021 0.091 0.066 0.155 0.011 0.063 0.362 0.450 0.681 0.280 PatchTST(2023) 0.185 0.298 0.022 0.093 0.080 0.183 0.011 0.066 0.136 0.266 0.271 0.122 SAITS(2023)1 0.196 0.289 0.215 0.190 0.087 0.163 0.135 0.139 0.445 0.512 0.932 0.484 GPT4TS(2024) 0.196 0.290 0.025 0.092 0.078 0.161 0.012 0.063 0.296 0.401 0.939 0.321 NuwaTS(specific) 0.182 0.293 0.020 0.091 0.070 0.164 0.011 0.064 0.086 0.191 0.307 0.127 PatchTST(one-for-all) 0.178 0.288 0.019 0.088 0.075 0.172 0.011 0.065 0.121 0.243 0.230 0.116 NuwaTS(one-for-all) 0.164 0.263 0.018 0.084 0.064 0.147 0.010 0.060 0.085 0.186 0.206 0.088 NuwaTS(fine-tuned) 0.156 0.255 0.017 0.082 0.060 0.142 0.010 0.058 0.081 0.183 0.207 0.086
@@
    We replace the MAE loss function in SAITS [8] and BRITS [6] with MSE loss function for fair comparison.
Table 2: Imputation results on four PEMS datasets with same setting as Table 1.
@@ 4.1 Main Results
We present the average MAE and MSE for specific, one-for-all, and fine-tuned NuwaTS models, along with other baselines, across 10 different datasets in Table 1 and Table 5. Furthermore, we visualize the average MSE across all datasets for the NuwaTS models and several baselines in Figure 5. On nearly all datasets, the specific NuwaTS model outperforms other domain-specific models. The comparison with GPT4TS [44] is particularly noteworthy. GPT4TS also involves training on a PLM backbone. However, it not only underperforms compared to NuwaTS, but it is also worse than PatchTST. This underscores the importance of the patching technique [^4] GPT4TS employs a channel-dependent approach on imputation task. in helping the model understand missing data patterns. Moreover, we observed that the generalization ability of NuwaTS and PatchTST (one-for-all) was further augmented after training on the general fused dataset, thereby affirming the existence of a scaling law [13] within the realm of time series modeling. We visualized the imputation results in Appendix C.1.
4.2 Zero-shot Capability
We validated the zero-shot capability of our model by directly evaluating a trained model on target data without further training. TimesNet and GPT4TS, which use a channel-dependent approach, can only perform zero-shot across datasets with the same number of variables, specifically the four ETT datasets in our study. In contrast, NuwaTS and PatchTST, as channel-independent approach, support inference on datasets with different variable dimension. We trained them on LargeST and then evaluated them on ECL and Weather datasets. The results in Table 3 shows NuwaTS exhibits superior zero-shot capability, achieving better performance on all cases. The reason may lie in the use of specialized missing data embedding and statistical embedding. Additionally, the language knowledge embedded in the PLM has strong generalization capabilities, whereas the PatchTST model trained on LargeST does not possess such strong capabilities.
Table 3: Zero shot evaluation results. All methods were trained on one dataset and zero shot to the other. A lower MSE or MAE indicates a better imputation performance. Red: the best. /: model failed to work.
@@ 4.2 Zero-shot Capability
$LargeST$$\Rightarrow$ $Weather$ 0.217 0.087 0.236 0.121 / / / /
@@ 4.3 Few-shot Domain-Specific Fine-Tuning
The ability to generalize to the target domain with a small amount of data is an important criterion for evaluating the generality of a model. We use $10\%$ and $1\%$ of the data for fine-tuning the NuwaTS model trained on LargeST. The results shown in Table 4 indicate that domain-specific fine-tuning is particularly effective for fields with limited data. On the ETT datasets, we achieved the same results using only 10% of the data as we did using 100% of the data.
@@ 4.4 Ablation Study
Table 5: Ablation results of the model training on LargeST and zero-shot on ECL and Weather. A lower MSE or MAE indicates a better imputation performance. Red: the best.
[b] Model ECL Weather MSE MAE MSE MAE Default 0.338 0.366 0.217 0.087 w/o-StatisticEmbedding 0.371 0.383 0.224 0.093 w/o-ContrastiveLearning 0.348 0.366 0.218 0.087 w/o-MissingEmbedding 0.338 0.368 0.221 0.092 w/o-GPT2weight1 0.339 0.368 0.218 0.089 FrozenBackbone 0.533 0.560 0.290 0.131
@@
We conducted ablation experiments via zero-shot evaluation. We trained the default model, several ablated models on LargeST. We then performed zero-shot evaluation on ECL and Weather datasets for out-of-domain evaluation. As shown in Table 5, after ablation of the Statistic Embedding, Missing Embedding, and contrastive learning module, the model’s zero-shot ability decreases significantly, proving the necessity of these key components. Freezing the PLM backbone leads to poor performance. Finally, we discovered that when training the model from scratch without using weights pre-trained on NLP tasks, its zero-shot generalization ability on other datasets deteriorates significantly. This indicates that training the model on NLP tasks benefits its performance on time series tasks, demonstrating that cross-modality training is meaningful [42].
We also verified the design of domain-specific fine-tuning strategy of NuwaTS in Table 6. When the randomly initialized prefix key&value and the knowledge-transfer layer were ablated, the model’s fine-tuning performance showed a significant decline which indicates that domain-transfer layer retains the learned knowledge while newly added prefix key&value provide flexibility in transferring to specific domain. Additionally, we find that merely appending the prefix to the input layer and fine-tuning it significantly reduces the effectiveness of fine-tuning. Inserting the prefix into each layer of the PLM enhances the prefix’s representation capacity during domain transfer [17].
Table 6: Domain-specific fine-tuning ablation study on ETTs based on the model pre-trained on general fused dataset. A lower MSE or MAE indicates a better imputation performance. Red: the best.
Model ETTh1 ETTh2 ETTm1 ETTm2
MSE MAE MSE MAE MSE MAE MSE MAE
no fine-tuning 0.164 0.263 0.018 0.084 0.064 0.147 0.010 0.060

### v3

@@ 3.3 Domain-Specific Fine-Tuning
To achieve a domain-specific model, we borrow the idea of P-tuningv2 strategy (Liu et al., 2022). Unlike pre-training phase, where the domain-specific embedding $\mathbf{k}$ is only added to the input embedding.
We incorporate this domain-specific information into every layer of the PLM. Initially, we employ a domain-transfer layer—a two-layer MLP network—to map $\mathbf{k}\in\mathbb{R}^{D}$ to $\bm{\mathcal{\hat{K}}}\in\mathbb{R}^{2\times\text{Layer}\times D}$.
@@
Following the P-tuningv2 approach, we also randomly initialize a continuous prompt $\bm{\mathcal{P}}\in\mathbb{R}^{2\times\text{Layer}\times D}$ for each layer. We then combine the randomly initialized $\bm{\mathcal{P}}$ with the domain-specific information $\bm{\mathcal{\hat{K}}}$ to serve as the domain-specific prefix Key and Value in the PLM’s hidden layers. Thus, we have $[\mathbf{Key}_{p},\mathbf{Value}_{p}]=\bm{\mathcal{P}}+\beta\bm{\mathcal{\hat{K}}}$, where $\beta=0.01$. Here, $\mathbf{Key}_{p}$ and $\mathbf{Value}_{p}\in\mathbb{R}^{\text{Layer}\times D}$ contain the domain-specific prefix key and value for every layer of the pre-trained model, thereby enhancing its representational capacity.
During fine-tuning on time series data from the target domain, we freeze all parameters except for the randomly initialized $\bm{\mathcal{P}}$ and the domain-transfer layer. The fine-tuning process is illustrated in Figure 3. If the domain-specific prefix is removed, the domain-specific model reverts to the original one-for-all foundation model. Therefore, the fine-tuning strategy we propose is essentially “plug-and-play”. Additionally, the domain-specific prefix is lightweight. For example, when using GPT-2 as the PLM backbone, the prefix requires less than 100KB of storage, compared to the 331.77MB required for the entire model. This makes our approach highly practical for real-world deployment, particularly in edge computing environments, as a single NuwaTS model can be trained on large-scale time series data. When deploying domain-specific models locally, only the corresponding prefix needs to be stored, significantly reducing storage requirements and enabling efficient deployment with minimal computational resources.
Additionally, we can incorporate inter-series correlations into the prefix. For the subsequent forecasting tasks in section 4.7, we design an inter-variable fine-tuning network to generate the layer prefix, which contains inter-series correlation information. Specific implementation details can be found in Algorithm 3 and Section B.4.
@@ 4 Experiment
We conducted comprehensive experiments on NuwaTS across ten commonly used datasets from various domains. To facilitate comparison, we provided four versions of the model based on the GPT2 architecture in Figure 4: (a) a specific model trained on a single domain, (b) an one-for-all model trained on a general fused dataset (17.6M collected samples from various domains), (c) a fine-tuned model for a specific domain (fine-tuned based on the one-for-all model) and (d) a cross-domain model pretrained only on LargeST (Liu et al., 2024b) (comprising 100.1 million collected samples). Cross-domain model will be evaluated on time series from other domains to verify the domain-transfer zero-shot capability in Section 4.3. Moreover, we evaluated BERT (Devlin et al., 2019) and LLaMA2 (Touvron et al., 2023) as backbones for NuwaTS (details in Appendix D.5).
Figure 4: The four different versions of NuwaTS trained in this study.
@@
Baselines. Following the configuration outlined in (Du et al., 2023), we evaluated two naive imputation methods: Median, where missing values are filled with the median value, and Last, where missing values are filled with the last previous observations. To ensure a fair comparison, all methods followed a unified pipeline [^1] https://github.com/thuml/Time-Series-Library. We assessed two classic deep learning-based imputation models, BRITS(Cao et al., 2018) and SAITS(Du et al., 2023). We also compared our model with other foundation models for time series such as DLinear(Zeng et al., 2023), PatchTST(Nie et al., 2023), iTransformer(Liu et al., 2024c), TimesNet(Wu et al., 2022), Autoformer(Wu et al., 2021), Fedformer(Zhou et al., 2022), and GPT4TS(Zhou et al., 2024). We trained GPT4TS separately on single datasets with the number of layers set to 6 to align with NuwaTS. We trained all these models using a domain-specific approach. Since PatchTST is a transformer-based model with bi-directional attention, we also trained and tested it on the general fused dataset.
@@
Setups. We partitioned the dataset along the sensor (variable) dimension into training, validation, and test sets in a 1:1:1 ratio. This division simulates the process of collecting relatively complete time series data from a few sensors or variables with lower missing rates, training a model with this data, and then using it to impute data from other sensors or variables that have higher missing rates. The input time series length $L$ is 96, and we trained all baselines under random missing rates (sampled from 0.1 to 0.9) and tested them separately under 9 missing rates: 0.1, 0.2,…, 0.9. We used a total of ten datasets for domain-specific training and mixed them into a general fused dataset for one-for-all training. Appendix A.1 shows the details of datasets. For the cross-domain model, we chose LargeST [^2] https://github.com/liuxu77/LargeST, a large-scale traffic dataset including a total of 8,600 time series and over 100 million samples, to train NuwaTS and conduct zero-shot experiments on other domains such as ETTs, weather and electricity.
@@
Table 1: Imputation results are presented for ETTs, ECL, and weather datasets with nine test missing rates: 0.1, 0.2, …, 0.9. All results are averaged across the nine different missing rates. A lower MSE or MAE signifies better imputation performance.  Green  highlights the best results, while  Yellow  indicates the second-best.
Model ETTh1 ETTh2 ETTm1 ETTm2 ECL Weather
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Median 0.723 0.611 0.728 0.472 0.699 0.584 0.744 0.464 1.010 0.834 0.998 0.497
Last 0.432 0.476 0.089 0.149 0.336 0.413 0.059 0.122 0.965 0.826 0.731 0.349
Autoformer(2021) 0.552 0.547 0.472 0.420 0.346 0.409 0.211 0.312 0.137 0.262 0.610 0.450
Fedformer(2022) 0.354 0.436 0.538 0.429 0.076 0.185 0.055 0.156 0.129 0.258 0.237 0.211
Dlinear(2023) 0.356 0.414 0.245 0.301 0.274 0.357 0.266 0.351 0.317 0.419 0.355 0.309
iTransformer(2024) 0.639 0.572 0.369 0.376 0.352 0.413 0.356 0.416 0.092 0.200 0.515 0.412
BRITS(2018) 0.213 0.313 0.117 0.172 0.096 0.183 0.071 0.123 0.317 0.427 0.793 0.474
TimesNet(2022) 0.166 0.280 0.021 0.091 0.066 0.155 0.011 0.063 0.362 0.450 0.681 0.280
PatchTST(2023) 0.185 0.298 0.022 0.093 0.080 0.183 0.011 0.066 0.136 0.266 0.271 0.122
SAITS(2023)1 0.196 0.289 0.215 0.190 0.087 0.163 0.135 0.139 0.445 0.512 0.932 0.484
GPT4TS(2024) 0.196 0.290 0.025 0.092 0.078 0.161 0.012 0.063 0.296 0.401 0.939 0.321
NuwaTS(specific) 0.182 0.293 0.020 0.091 0.070 0.164 0.011 0.064 0.086 0.191 0.307 0.127
PatchTST(one-for-all) 0.178 0.288 0.019 0.088 0.075 0.172 0.011 0.065 0.121 0.243 0.230 0.116
NuwaTS(one-for-all)  0.164  0.263  0.018  0.084  0.064  0.147  0.010  0.060  0.085  0.186  0.206  0.088
NuwaTS(fine-tuned)  0.156  0.255  0.017  0.082  0.060  0.142  0.010  0.058  0.081  0.183  0.207  0.086
@@
    We replace the MAE loss function in SAITS (Du et al., 2023) and BRITS (Cao et al., 2018) with MSE loss function for fair comparison.
Table 2: Imputation results on four PEMS datasets with the same setting as Table 1.
@@ 4.1 Main Results
We present the average MAE and MSE across different missing rates for specific, one-for-all, and fine-tuned NuwaTS models, along with other baselines, across 10 different datasets in Table 1 and Table 2, detailed results are shown in Appendix E. Furthermore, we visualize the average MSE across all datasets at various missing rates for the NuwaTS models and several baselines, as shown in Figure 5. On nearly all datasets, the NuwaTS model outperforms other domain-specific models. Moreover, we observed that the generalization ability of both NuwaTS and PatchTST (one-for-all) was further enhanced after training on a fused dataset spanning multiple domains, supporting the presence of a scaling law (Kaplan et al., 2020) in time series imputation tasks. We visualized the imputation results in Appendix C.1. We also conducted experiments to verify NuwaTS on real-world dataset in Appendix D.1.
4.2 Imputation Results under Continuous Missing
In Section 4.1, we conduct a comprehensive evaluation of the performance of different approaches under purely random missing data scenarios. In real-world scenarios, missing data often occurs in a continuous manner (Du et al., 2024). Therefore, we also evaluated NuwaTS’s performance under conditions of continuous missing data. The experimental setup is set as the following: Assuming a missing rate of $r$, we randomly selected a continuous segment of the time series with a length of $L\times r$ for masking. We evaluate the models trained on randomly missing data by directly testing them on datasets with continuous missing patterns.
Table 3: Imputation results on continuous missing data. The final results are averaged across 9 missing rates: 0.1, 0.2… 0.9. A lower MSE or MAE indicates better imputation performance.  Green : the best.
Model ETTh1 ETTh2
MSE MAE MSE MAE
BRITS 0.717 0.619 0.580 0.519
SAITS 0.519 0.469 0.318 0.250
TimesNet 0.497 0.494 0.110 0.154
GPT4TS 0.483 0.480 0.104 0.154
NuwaTS  0.465  0.455  0.095  0.141
As it is shown in Table 3, our findings reveal that NuwaTS exhibits strong domain adaptation capabilities and is highly robust in handling diverse missing patterns, further highlighting its generalizability across varying data missing patterns.
4.3 Domain-Transfer Zero-shot Capability
We validated the zero-shot capability of our model by directly evaluating a trained model on target data without further training. TimesNet and GPT4TS, which use a channel-dependent approach, can only perform zero-shot across datasets with the same number of variables, specifically the four ETT datasets in our study. In contrast, NuwaTS and PatchTST, as channel-independent approaches, support inference on datasets with different variable dimensions. We trained them on LargeST and then evaluated them on ECL and Weather datasets. The results in Table 4 shows NuwaTS exhibits superior zero-shot capability, achieving better performance on all cases. The reason may lie in the use of specialized missing data embedding and statistical embedding. Additionally, the language knowledge embedded in the PLM has strong generalization capabilities, whereas the PatchTST model trained on LargeST does not possess such strong capabilities.
Table 4: Zero shot evaluation results. All methods were trained on one dataset and zero shot to the other. A lower MSE or MAE indicates better imputation performance.  Green : the best. /: model failed to work.
@@ 4.3 Domain-Transfer Zero-shot Capability
$LargeST$Rightarrow $Weather$  0.217  0.087 0.236 0.121 / / / /
@@ 4.4 Few-shot Domain-Specific Fine-Tuning
The ability to generalize to the target domain with a small amount of data is an important criterion for evaluating the generality of a model. We use $10\%$ and $1\%$ of the data for fine-tuning the cross-domain NuwaTS model which is trained on LargeST. The results shown in Table 5 indicate that domain-specific fine-tuning is particularly effective for fields with limited data. On the ETT datasets, we achieved the same results using only 10% of the data as we did using 100% of the data.
@@ 4.5 Ablation Study
Table 6: Ablation results of the model training on ETTh1 and LargeST and zero-shot on other data domain. A lower MSE or MAE indicates better imputation performance.  Green : the best.
[b] Model ETTh1 $\Rightarrow$ ETTh2 ETTh1 $\Rightarrow$ Weather LargeST $\Rightarrow$ ECL LargeST $\Rightarrow$ Weather MSE MAE MSE MAE MSE MAE MSE MAE Default  0.023  0.091  0.284  0.164  0.338  0.366  0.217  0.087 w/o-StatisticEmbedding 0.024 0.091 0.303 0.176 0.371 0.383 0.224 0.093 w/o-ContrastiveLearning 0.025 0.097 0.311 0.184 0.348 0.366 0.218 0.087 w/o-MissingEmbedding 0.028 0.101 0.331 0.193 0.338 0.368 0.221 0.092 w/o-GPT2weight1 0.024 0.093 0.311 0.177 0.339 0.368 0.218 0.089 FrozenBackbone 0.027 0.095 0.344 0.207 0.533 0.560 0.290 0.131
@@
We conducted ablation experiments via zero-shot evaluation. We trained the default model, several ablated models on ETTh1 with 34650 samples and LargeST with more than 100.1 million samples. We then performed zero-shot evaluation on other datasets for out-of-domain evaluation. As shown in Table 6, after ablation of the Statistic Embedding, Missing Embedding, and contrastive learning module, the model’s zero-shot ability decreases significantly, proving the necessity of these key components. Freezing the PLM backbone leads to poor performance. Finally, we discovered that when training the model from scratch without using weights pre-trained on NLP tasks, its zero-shot generalization ability on other datasets deteriorates significantly. This indicates that training the model on NLP tasks benefits its performance on time series tasks, demonstrating that cross-modality training is meaningful (Zhang et al., 2024).
When comparing the ablation results on ETTh1 and LargeST, where the former is a very small dataset and the latter is significantly larger. Our findings show that the specially designed tokens and contrastive learning modules provide more noticeable improvements on the smaller ETTh1 dataset. This suggests that for foundation models, data quantity might be more critical than the inductive biases introduced by specialized modules.